# Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

### v1

@@ 4.2 Convergence with High Probability and Markovian Samples
By applying the $L^{p}$ convergence result (Theorem D.1) with $p=2\log(1/\delta)$ and Markov’s inequality, we obtain the high-probability convergence result.
@@
For any $\varepsilon>0$ and $\delta\in(0,1)$, suppose $K\geq(1{-}\gamma)^{-1}$, $\alpha\in(0,(1{-}\sqrt{\gamma}){/}[38\log(T{/}\delta^{2})])$, ${\bm{w}}_{0}\in{\mathbb{R}}^{dK}$ is the initialization, and total update steps $T=$
$\widetilde{O}\Bigg{(}$ $\frac{\frac{1}{K(1{-}\gamma)^{2}}\!\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}{\otimes}{\bm{\Sigma}}_{{\bm{\phi}}}}\!\!\!{+}1}{\varepsilon^{2}(1-\gamma)^{2}\lambda_{\min}}\!\left({1{+}\frac{\alpha\log\frac{1}{\delta}}{(1{-}\gamma)\lambda_{\min}}}\right)\!\log\!\frac{1}{\delta}{+}\frac{\frac{1}{\sqrt{K}(1{-}\gamma)}\!\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}{\otimes}{\bm{\Sigma}}_{{\bm{\phi}}}}\!\!\!{+}1}{\varepsilon\sqrt{\alpha}(1-\gamma)^{\frac{3}{2}}\lambda_{\min}}\!\log\!\frac{1}{\delta}{+}\frac{\log\!\frac{\left\|{{\bm{\theta}}_{0}{-}{\bm{\theta}}^{\star}}\right\|}{\varepsilon\sqrt{K}(1{-}\gamma)}}{\alpha(1{-}\gamma)\lambda_{\min}}\Bigg{)},$
then with probability at least $1-\delta$, it holds that ${\mathcal{L}}\left({\bar{\bm{w}}_{T}}\right)\leq\varepsilon$. Here, the $\widetilde{O}\left({\cdot}\right)$ does not hide polynomials of $\log(1/\delta)$ (but hides logarithm terms of $\log(1/\delta)$).
Again, we will obtain concrete sample complexity bounds as in Eqn. (10) or Eqn. (11) if we use different step sizes, we omit these for brevity. Compared with the theoretical results for classic Linear-TD, our results match (Theorem 4 Samsonov et al., 2024b) that also considers the constant step size, but has a worse dependence on $\log\left({1/\delta}\right)$ than (Theorem 3.1 Wu et al., 2024) which considers the polynomial-decaying step size $\alpha_{t}=\alpha_{0}t^{-\beta}$ with $\beta\in(1/2,1)$ instead.
@@
Using the same argument as in the proof of (Theorem 6 Samsonov et al., 2024b), one can immediately derive a high-probability sample complexity bound in the Markovian setting. Compared with the bound in the generative model setting (Theorem 4.2), the bound in the Markovian setting will have an additional dependency on $t_{\operatorname{mix}}\log(T/\delta)$, where $t_{\operatorname{mix}}$ is the mixing time of the Markov chain $\left\{{s_{t}}\right\}_{t=0}^{\infty}$ in ${\mathcal{S}}$. We omit this result for brevity.
@@ 5.1 Vectorization of Linear-CTD
In our analysis, it will be more convenient to work with the vectorization version and the CDF representation introduced in Appendix C.6. In Appendix C.7, we show that Linear-CTD (Eqn. (9)) is equivalent to SSGD with the CDF representation (Eqn. (24)). We consider the vectorization of Eqn. (24), and we denote ${\bm{\theta}}_{t}=\left({{\bm{C}}\otimes{\bm{I}}_{d}}\right){\bm{w}}_{t}$ (other CDF parameters are defined in the same way):
${\bm{\theta}}_{t}{=}{\bm{\theta}}_{t-1}{-}\alpha\left({{\bm{A}}_{t}{\bm{\theta}}_{t-1}{-}{\bm{b}}_{t}}\right),\quad{\bm{A}}_{t}{=}\left[{{\bm{I}}_{K}{\otimes}\left({{\bm{\phi}}(s_{t}){\bm{\phi}}(s_{t})^{\intercal}}\right)}\right]{-}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r_{t}){\bm{C}}^{-1}}\right){\otimes}\left({{\bm{\phi}}(s_{t}){\bm{\phi}}(s^{\prime}_{t})^{\intercal}}\right)}\right],$
${\bm{b}}_{t}=\frac{1}{K+1}\left[{{\bm{C}}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)}\right]\otimes{\bm{\phi}}(s_{t}).$ (13)
We denote by $\bar{{\bm{A}}},\bar{{\bm{b}}}$ the expectation of ${\bm{A}}_{t},{\bm{b}}_{t}$ respectively. Utilizing the exponential stability arguments, we can derive an upper bound for $\left\|{\bar{{\bm{A}}}\left({\bar{{\bm{\theta}}}_{T}{-}{\bm{\theta}}^{\star}}\right)}\right\|$. We need to further translate it to an upper bound for ${\mathcal{L}}(\bar{{\bm{w}}}_{T})$, which is done in the following lemma, whose proof can be found in Appendix D.1.
@@
For any ${\bm{w}}\in{\mathbb{R}}^{dK}$, it holds that ${\mathcal{L}}({\bm{w}})\leq 2{K^{-1/2}(1{-}\gamma)^{-2}\lambda_{\min}^{-1/2}}\left\|{\bar{{\bm{A}}}\left({{{\bm{\theta}}}{-}{\bm{\theta}}^{\star}}\right)}\right\|$.
@@ 5.2 Exponential Stability Analysis
First, we introduce some notations. Let ${\bm{e}}_{t}:{=}{{\bm{A}}}_{t}{\bm{\theta}}^{\star}{-}{{\bm{b}}}_{t}$, we denote by $C_{A}$ (resp. $C_{e}$) the almost sure upper bound for $\max\left\{{\left\|{{\bm{A}}_{t}}\right\|,\left\|{{\bm{A}}_{t}{-}\bar{{\bm{A}}}}\right\|}\right\}$ (resp. $\left\|{{\bm{e}}_{t}}\right\|$), and ${\bm{\Sigma}}_{e}:={\mathbb{E}}\left[{{\bm{e}}_{t}{\bm{e}}_{t}^{\intercal}}\right]$ the covariance matrix of ${\bm{e}}_{t}$. The following lemma provides useful upper bounds, whose proof can be found in Appendix D.2.
@@
Suppose $K\geq(1-\gamma)^{-1}$, then it holds that
@@
Let ${\bm{\Gamma}}_{m:n}^{(\alpha)}:=\prod_{i=m}^{n}\left({{\bm{I}}-\alpha{\bm{A}}_{i}}\right)$ for any $\alpha>0$ and $m,n\in{\mathbb{N}}$, $m\leq n$. The exponential stability of Linear-CTD is summarized in the following lemma, whose proof can be found in Appendix D.3.
@@
For any $p\geq 2$, let $a=(1-\sqrt{\gamma})\lambda_{\min}/2$ and $\alpha_{p,\infty}=(1-\sqrt{\gamma})/(38p)$ ($\alpha_{p,\infty}p\leq 1/2$). Then for any $\alpha\in\left({0,\alpha_{p,\infty}}\right)$, ${\bm{u}}\in{\mathbb{R}}^{dK}$ and $n\in{\mathbb{N}}$, it holds that ${\mathbb{E}}^{1/p}[\|{\bm{\Gamma}}_{1:n}^{(\alpha)}{\bm{u}}\|^{p}]\leq(1-\alpha a)^{n}\left\|{{\bm{u}}}\right\|$.
See Appendix D.6 for the counterparts of Lemma 5.1, Lemma 5.2 and Lemma 5.3 for vanilla SSGD with the PMF representation, which have additional dependency on $K$.
Combining Lemma 5.1, Lemma 5.2 and Lemma 5.3 with (Theorem 1 Samsonov et al., 2024b), one can immediately obtain Theorem 4.1, the remaining details can be found in Appendix D.4.
@@ 5.3 Key Steps in the Proofs
In this section, we highlight some key steps in proving above theoretical results.
Bounding Spectral Norm of Expectation of Kronecker Product.
In proving that the ${\mathcal{L}}({\bm{w}})$ can be upper-bounded by $\left\|{\bar{{\bm{A}}}\left({{{\bm{\theta}}}-{\bm{\theta}}^{\star}}\right)}\right\|$ (Lemma 5.1), as well as in verifying the exponential stability condition (Lemma 5.3), one of the most critical steps is to show
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|\leq\sqrt{\gamma},\quad\forall r\in[0,1].$ (14)
By Lemma E.3, we have $\|{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{{-}1}\|{\leq}\sqrt{\gamma}$ for any $r\in[0,1]$. In addition, one can check that $\|{\mathbb{E}}_{s,s^{\prime}}[{\bm{\Sigma}}_{{\bm{\phi}}}^{-1{/}2}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-1{/}2}]\|{\leq}1$. One may speculate that the property $\left\|{{\bm{B}}_{1}{\otimes}{\bm{B}}_{2}}\right\|=\left\|{{\bm{B}}_{1}}\right\|\left\|{{\bm{B}}_{2}}\right\|$ (Lemma A.3) is enough to get the desired conclusion. However, the two matrices in the Kronecker product are not independent, preventing us from using this simple property to derive the conclusion. On the other hand, since we merely have the upper bound ${\mathbb{E}}_{s,s^{\prime}}[\|{\bm{\Sigma}}_{{\bm{\phi}}}^{-1{/}2}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-1{/}2}\|]{\leq}d$, simply moving the expectation in Eqn. (14) outside the norm will lead to a loose $d\sqrt{\gamma}$ bound . To resolve this problem, we leverage the fact that the second matrix is rank-$1$ and prove the following result. The proof can be found in the derivation following Eqn. (26).
@@
For any random matrix ${\bm{Y}}$ and random vectors ${\bm{x}}$, ${\bm{z}}$, suppose $\left\|{{\bm{Y}}}\right\|\leq C_{Y}$ almost surely, ${\mathbb{E}}\left[{{\bm{x}}{\bm{x}}^{\intercal}}\right]\preccurlyeq C_{x}{\bm{I}}_{d_{1}}$ and ${\mathbb{E}}\left[{{\bm{z}}{\bm{z}}^{\intercal}}\right]\preccurlyeq C_{z}{\bm{I}}_{d_{2}}$ for some constants $C_{Y},C_{x},C_{z}>0$, then it holds that
$\left\|{{\mathbb{E}}\left[{{\bm{Y}}\otimes\left({{\bm{x}}{\bm{z}}^{\intercal}}\right)}\right]}\right\|\leq C_{Y}\sqrt{C_{x}C_{z}}.$
Remark 4.
The matrix ${\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}$ also appeared in (Proposition B.2 Rowland et al., 2024b) as the matrix representation of the categorical projected Bellman operator ${\bm{\Pi}}_{K}\bm{\mathcal{T}}^{\pi}$ of an one-state MDP. They showed $\|{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{{-}1}\|{\leq}\sqrt{\gamma}$ using the fact that ${\bm{\Pi}}_{K}\bm{\mathcal{T}}^{\pi}$ is a $\sqrt{\gamma}$-contraction in $\left({{\mathscr{P}},\ell_{2}}\right)$. Our Lemma E.3 provides a new analysis by directly analyzing the matrix. See Appendix F for more details.
Bounding Norm of ${\bm{b}}_{t}$.
In proving Lemma 5.2, the most involved step is to upper-bound $\left\|{{\bm{b}}_{t}}\right\|$ (Eqn. (13)). To this end, we need to upper-bound the following term which appears in Eqn. (27):
$(K+1)^{-1}\|\sum_{j=0}^{K}{\bm{C}}{\bm{g}}_{j}(r)-{\bm{C}}\bm{1}_{K}\|,\quad\forall r\in[0,1],$ (15)
when $K\geq(1-\gamma)^{-1}$. Same as the matrix ${\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}$ (see Remark 4), Term (15) is also related to the matrix representation of the categorical projected Bellman operator. Specifically, let $\nu=(K+1)^{-1}\sum_{k=0}^{K}\delta_{x_{k}}$ be the discrete uniform distribution, Term (15) equals
@@
where we used the orthogonal decomposition (Proposition B.2) and an upper bound for $\ell_{2}\left({(b_{r,\gamma})_{\#}(\nu),\nu}\right)$ (Lemma E.4). The full proof can be found in the derivation following Eqn. (27).

### v2

@@ 4.2 Convergence with High Probability and Markovian Samples
Applying the $L^{p}$ convergence result (Theorem E.1) with $p=2\log(1/\delta)$ and Markov’s inequality, we immediately obtain the high-probability convergence result.
@@
For any $\varepsilon>0$ and $\delta\in(0,1)$, suppose $K\geq(1-\gamma)^{-1}$, $\alpha\in(0,(1-\sqrt{\gamma}){/}[38\log(T/\delta^{2})])$, and
$T=\widetilde{{\mathcal{O}}}\Bigg{(}$ $\frac{\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{V_{1}}{+}1}{\varepsilon^{2}(1-\gamma)^{2}\lambda_{\min}}\left({1{+}\frac{\alpha\log\frac{1}{\delta}}{(1{-}\gamma)\lambda_{\min}}}\right)\log\frac{1}{\delta}{+}\frac{\left\|{{\bm{\theta}}^{\star}}\right\|_{V_{1}}{+}1}{\varepsilon\sqrt{\alpha}(1-\gamma)^{\frac{3}{2}}\lambda_{\min}}\log\frac{1}{\delta}{+}\frac{\log\frac{\left\|{{\bm{\theta}}_{0}{-}{\bm{\theta}}^{\star}}\right\|_{V_{2}}}{\varepsilon}}{\alpha(1{-}\gamma)\lambda_{\min}}\Bigg{)}.$
Then with probability at least $1-\delta$, it holds that ${\mathcal{L}}\left({\bar{\bm{\theta}}_{T}}\right)\leq\varepsilon$. Here, the $\widetilde{O}\left({\cdot}\right)$ does not hide polynomials of $\log(1/\delta)$ (but hides logarithm terms of $\log(1/\delta)$).
Again, we will obtain concrete sample complexity bounds as in Eqn. (14) or Eqn. (15) if we use different step sizes. Compared with the theoretical results for classic Linear-TD, our results match Samsonov et al. [54, Theorem 4]. [54] also considered the constant step size, but obtained a worse dependence on $\log\left({1/\delta}\right)$ than Wu et al. [69, Theorem 4] which uses the polynomial-decaying step size $\alpha_{t}=\alpha_{0}t^{-\beta}$ with $\beta\in(1/2,1)$ instead.
@@
Using the same argument as in the proof of Samsonov et al. [54, Theorem 6], one can immediately derive a high-probability sample complexity bound in the Markovian setting. Compared with the bound in the generative model setting (Theorem 4.2), the bound in the Markovian setting will have an additional dependency on $t_{\operatorname{mix}}\log(T/\delta)$, where $t_{\operatorname{mix}}$ is the mixing time of the Markov chain $\left\{{s_{t}}\right\}_{t=0}^{\infty}$ in ${\mathcal{S}}$. We omit this result for brevity.
@@ 5.1 Vectorization of Linear-CTD
In our analysis, it will be more convenient to work with the vectorization version of the updating scheme of Linear-CTD (Eqn. (13)):
${\bm{\theta}}_{t}{\leftarrow}{\bm{\theta}}_{t-1}{-}\alpha\left({{\bm{A}}_{t}{\bm{\theta}}_{t-1}{-}{\bm{b}}_{t}}\right),\quad{\bm{A}}_{t}{=}\left[{{\bm{I}}_{K}{\otimes}\left({{\bm{\phi}}(s_{t}){\bm{\phi}}(s_{t})^{\top}}\right)}\right]{-}[({\bm{C}}\tilde{{\bm{G}}}(r_{t}){\bm{C}}^{-1}){\otimes}\left({{\bm{\phi}}(s_{t}){\bm{\phi}}(s^{\prime}_{t})^{\top}}\right)],$
${\bm{b}}_{t}=(K+1)^{-1}\left[{{\bm{C}}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)}\right]\otimes{\bm{\phi}}(s_{t}).$ (16)
We denote by $\bar{{\bm{A}}}$ and $\bar{{\bm{b}}}$ the expectations of ${\bm{A}}_{t}$ and ${\bm{b}}_{t}$, respectively. Using exponential stability arguments, we can derive an upper bound for $\|\bar{{\bm{A}}}(\bar{{\bm{\theta}}}_{T}-{\bm{\theta}}^{\star})\|$. The following lemma further translates it to an upper bound for ${\mathcal{L}}(\bar{{\bm{\theta}}}_{T})=W_{1,\mu_{\pi}}(\bm{\eta}_{{\bm{\theta}}^{\star}},\bm{\eta}_{\bar{\bm{\theta}}_{T}})$, whose proof is given in Appendix E.1.
@@
For any ${\bm{\theta}}\in{\mathbb{R}}^{dK}$, it holds that ${\mathcal{L}}({\bm{\theta}})\leq 2{K^{-1{/}2}(1-\gamma)^{-2}\lambda_{\min}^{-1{/}2}}\left\|{\bar{{\bm{A}}}\left({{{\bm{\theta}}}-{\bm{\theta}}^{\star}}\right)}\right\|$.
@@ 5.2 Exponential Stability Analysis
First, we introduce some notations. Letting ${\bm{e}}_{t}:{=}{{\bm{A}}}_{t}{\bm{\theta}}^{\star}{-}{{\bm{b}}}_{t}$, we denote by $C_{A}$ (resp. $C_{e}$) the almost sure upper bound for $\max\{\|{\bm{A}}_{t}\|,\|{\bm{A}}_{t}{-}\bar{{\bm{A}}}\|\}$ (resp. $\|{\bm{e}}_{t}\|$), and ${\bm{\Sigma}}_{e}:={\mathbb{E}}[{\bm{e}}_{t}{\bm{e}}_{t}^{\top}]$ the covariance matrix of ${\bm{e}}_{t}$. The following lemma provides useful upper bounds, whose proof is given in Appendix E.2.
@@
For any $K\geq(1-\gamma)^{-1}$, it holds that
@@
Let ${\bm{\Gamma}}_{t}^{(\alpha)}:=\prod_{i=1}^{t}({\bm{I}}-\alpha{\bm{A}}_{i})$ for any $\alpha>0$ and $t\in{\mathbb{N}}$. The exponential stability of Linear-CTD is summarized in the following lemma, whose proof can be found in Appendix E.3.
@@
For any $p\geq 2$, let $a=(1-\sqrt{\gamma})\lambda_{\min}/2$ and $\alpha_{p,\infty}=(1-\sqrt{\gamma})/(38p)$ ($\alpha_{p,\infty}p\leq 1/2$). Then for any $\alpha\in\left({0,\alpha_{p,\infty}}\right)$, ${\bm{u}}\in{\mathbb{R}}^{dK}$ and $t\in{\mathbb{N}}$, it holds that ${\mathbb{E}}^{1/p}[\|{\bm{\Gamma}}_{t}^{(\alpha)}{\bm{u}}\|^{p}]\leq(1-\alpha a)^{t}\left\|{{\bm{u}}}\right\|$.
The following theorem states the $L^{2}$ convergence of $\|\bar{{\bm{A}}}(\bar{{\bm{\theta}}}_{T}{-}{\bm{\theta}}^{\star})\|$ based on exponential stability arguments. For the general $L^{p}$ convergence, please refer to Samsonov et al. [54, Theorem 2].
Theorem 5.1.
[54, Theorem 1] For any $\alpha\in(0,\alpha_{2,\infty})$, it holds that
${\mathbb{E}}^{1/2}[\|\bar{{\bm{A}}}(\bar{{\bm{\theta}}}_{T}{-}{\bm{\theta}}^{\star})\|^{2}]\lesssim$ $\frac{\sqrt{\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)}}{\sqrt{T}}\!\left({1{+}\frac{C_{A}\sqrt{\alpha}}{\sqrt{a}}}\right)+\frac{\sqrt{\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)}}{T\sqrt{\alpha a}}+\frac{(1{-}\alpha a)^{T{/}2}}{T}\!\left({\frac{1}{\alpha}{+}\frac{C_{A}}{\sqrt{\alpha a}}}\right)\left\|{{\bm{\theta}}_{0}{-}{\bm{\theta}}^{\star}}\right\|.$
Combining these lemmas with Theorem 5.1, we can immediately obtain Theorem 4.1.
Remark 5 (Convergence of SSGD with PMF representation).
In Appendix E.5, we give counterparts of these lemmas and Theorem 4.1 for vanilla SSGD with PMF representation. The results imply that the step size in the algorithm should scale with $(1{-}\sqrt{\gamma}){/}K^{2}$ and the sample complexity grows with $K$. In Appendix F.2, we verify through numerical experiments that as $K$ increases, to ensure convergence, the step size of the vanilla algorithm indeed needs to decay at a rate of $K^{-2}$. In contrast, the step size of our Linear-CTD does not need to be adjusted when $K$ increases. Moreover, we find that Linear-CTD empirically consistently outperforms the vanilla algorithm under different $K$.
@@ 5.3 Key Steps in the Proofs
Here we highlight some key steps in proving the above theoretical results.
Bounding the Spectral Norm of Expectation of Kronecker Products.
In proving that the ${\mathcal{L}}({\bm{\theta}})$ can be upper-bounded by $\|\bar{{\bm{A}}}({{\bm{\theta}}}-{\bm{\theta}}^{\star})\|$ (Lemma 5.1), as well as in verifying the exponential stability condition (Lemma 5.3), one of the most critical steps is to show
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|\leq\sqrt{\gamma},\quad\forall r\in[0,1].$ (17)
By Lemma G.3, we have $\|{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{{-}1}\|{\leq}\sqrt{\gamma}$ for any $r\in[0,1]$. In addition, one can check that $\|{\mathbb{E}}_{s,s^{\prime}}[{\bm{\Sigma}}_{{\bm{\phi}}}^{-1{/}2}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-1{/}2}]\|{\leq}1$. One may speculate that the property $\|{\bm{B}}_{1}{\otimes}{\bm{B}}_{2}\|=\|{\bm{B}}_{1}\|\|{\bm{B}}_{2}\|$ (Lemma A.3) is enough to get the desired conclusion. However, the two matrices in the Kronecker product are not independent, preventing us from using this simple property to derive the conclusion. On the other hand, since we only have the upper bound ${\mathbb{E}}_{s,s^{\prime}}[\|{\bm{\Sigma}}_{{\bm{\phi}}}^{-1{/}2}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-1{/}2}\|]{\leq}d$, simply moving the expectation in Eqn. (17) outside the norm will lead to a loose $d\sqrt{\gamma}$ bound. To resolve this problem, we leverage the fact that the second matrix is rank-$1$ and prove the following result. The proof can be found in the derivation following Eqn. (29).
@@
For any random matrix ${\bm{Y}}$ and random vectors ${\bm{x}}$, ${\bm{z}}$, suppose $\left\|{{\bm{Y}}}\right\|\leq C_{Y}$ almost surely, ${\mathbb{E}}\left[{{\bm{x}}{\bm{x}}^{\top}}\right]\preccurlyeq C_{x}{\bm{I}}_{d_{1}}$ and ${\mathbb{E}}\left[{{\bm{z}}{\bm{z}}^{\top}}\right]\preccurlyeq C_{z}{\bm{I}}_{d_{2}}$ for some constants $C_{Y},C_{x},C_{z}>0$. Then it holds that
$\left\|{{\mathbb{E}}\left[{{\bm{Y}}\otimes\left({{\bm{x}}{\bm{z}}^{\top}}\right)}\right]}\right\|\leq C_{Y}\sqrt{C_{x}C_{z}}.$
Remark 6 (Matrix Representation of Categorical Projected Bellman operator).
The matrix ${\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}$ also appears in Rowland et al. [51, Proposition B.2] as the matrix representation of the categorical projected Bellman operator ${\bm{\Pi}}_{K}\bm{\mathcal{T}}^{\pi}$ of a specific one-state MDP. As a result, $\|{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{{-}1}\|{\leq}\sqrt{\gamma}$ because ${\bm{\Pi}}_{K}\bm{\mathcal{T}}^{\pi}$ is a $\sqrt{\gamma}$-contraction in $\left({{\mathscr{P}},\ell_{2}}\right)$. Our Lemma G.3 provides a new analysis by directly analyzing the matrix.
Bounding the Norm of ${\bm{b}}_{t}$.
In proving Lemma 5.2, the most involved step is to upper-bound $\left\|{{\bm{b}}_{t}}\right\|$ (Eqn. (16)). To this end, we need to upper-bound the following term:
$\frac{1}{K+1}\left\|{{\bm{C}}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)}\right\|,\quad\forall r\in[0,1].$ (18)
Term (18) is also related to the categorical projected Bellman operator ${\bm{\Pi}}_{K}\bm{\mathcal{T}}^{\pi}$. Specifically, let $\nu{=}\frac{1}{K{+}1}\sum_{k=0}^{K}\delta_{x_{k}}$ be the discrete uniform distribution. One can show that Term (18) equals
@@
where we used the fact that $\Pi_{K}$ is non-expansive and an upper bound for $\ell_{2}\left({(b_{r,\gamma})_{\#}(\nu),\nu}\right)$ when $K\geq(1-\gamma)^{-1}$ (Lemma G.4). The full proof can be found in the derivation following Eqn. (30).